A firm offers terms of 1.4/10, net 60. a. What effective annual interest rate does the firm earn when a customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What effective annual interest rate does the firm earn if the terms are changed to 2.4/10, net 60, and the customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) c. What effective annual interest rate does the firm earn if the terms are changed to 1.4/10, net 75, and the customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) d. What effective annual interest rate does the firm earn if the terms are changed to 1.4/15, net 60, and the customer does not take the discount? (Use 365 days a year. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
a] EAR = (1+1.4/88.6)^(365/50)-1 = 12.13%
b] EAR = (1+2.4/87.6)^(365/50)-1 = 21.81%
c] EAR = (1+1.4/88.6)^(365/65)-1 = 9.20%
d] EAR = (1+1.4/88.6)^(365/45)-1 = 13.56%